Grade 12Physical Chemistry

50 ml of 0.1M NH4OH, 25 ml of 2 M NH4Cl were used to make a buffer.what is the pH if pkb is 4.8?

lokesh palingi
12 Years agoGrade 12

1 Answer

Sunil Kumar FP
12 Years ago
it is a basic buffer sopution.
the formula is ph=14-[pkb +log(salt/base)
ph=14-(4.8 +log 50/5)
ph= 8.2
(salt=nh4cl=25×2=50mm), (base=nh4oh=.1×50 millimole=5)

thanks and regrds
sunil kr
askIItian faculty